### v1

@@ Abstract
Large language models (LLMs) have shown remarkable potential in processing long sequences, yet efficiently serving these long-context models remains challenging due to the quadratic computational complexity of attention in the prefilling stage and the large memory footprint of the KV cache in the decoding stage. To address these issues, we introduce LServe, an efficient system that accelerates long-sequence LLM serving via hybrid sparse attention. This method unifies different hardware-friendly, structured sparsity patterns for both prefilling and decoding attention into a single framework, where computations on less important tokens are skipped block-wise. LServe demonstrates the compatibility of static and dynamic sparsity in long-context LLM attention. This design enables multiplicative speedups by combining these optimizations. Specifically, we convert half of the attention heads to nearly free streaming heads in both the prefilling and decoding stages. Additionally, we find that only a constant number of KV pages is required to preserve long-context capabilities, irrespective of context length. We then design a hierarchical KV page selection policy that dynamically prunes KV pages based on query-centric similarity. On average, LServe accelerates LLM prefilling by up to 2.9$\times$ and decoding by 1.3-2.1$\times$ over vLLM, maintaining long-context accuracy. Code is released at https://github.com/mit-han-lab/omniserve.
footnotetext: * indicates equal contribution. Part of the work was done while Shang Yang and Haotian Tang are interning at NVIDIA. 1MIT 2SJTU 3NVIDIA.
@@ 1 Introduction
Large Language Models (LLMs) have dramatically transformed the field of artificial intelligence. With expanding context window lengths, LLMs now demonstrate remarkable performance across diverse long-sequence applications Gemini Team, Google (2024), including multi-turn conversations, long document analysis Zhang et al. (2024b); Goyal & Durrett (2020); Huang et al. (2021), multi-modal understanding Xue et al. (2024a); Liu et al. (2024b); Lin et al. (2024a), and code completion Li et al. (2023); Lozhkov et al. (2024). Many of these applications require processing hundreds of thousands of context tokens in real-world settings, presenting unique challenges. In particular, the demand for fast prefilling, or minimizing the time to the first token, and the burden on the decoding phase due to the large KV (key-value) caches necessary for such contexts, represent significant hurdles.
@@
We benchmark LServe across three long-sequence LLMs—Llama-3-8B, Minitron-4B, and Llama-2-7B—at context lengths up to 512k tokens. Compared to state-of-the-art frameworks like vLLM Kwon et al. (2023b), QServe Lin et al. (2024b), MInference Jiang et al. (2024b), and DuoAttention Xiao et al. (2024), LServe accelerates prefilling stage by up to 2.9$\times$ and achieves an average of 1.3$\times$-2.1$\times$ speedup in the decoding stage. Furthermore, LServe accomplishes these speedups while retaining the long-context capabilities of the original dense, floating-point models, demonstrating that hybrid attention sparsity is a free lunch for long-sequence LLM serving.
@@ 2.2 Motivation
Figure 2: Latency breakdown of LLM inference for both prefilling and decoding stage. As sequence length increases, attention dominates both stages due to its quadratic complexity in prefilling stage and linear complexity during decoding stage. In contrast, GEMM exhibits linear complexity during prefilling stage and constant complexity during decoding stage. Latency numbers measured with Llama-3-8B on NVIDIA A100 GPU.
@@ 6 Related Work
Sparse Attention. BigBird Zaheer et al. (2020) reduces attention complexity by blending local, global, and random attention masks. Subsequent methods like StreamingLLM Xiao et al. (2023), H2O Zhang et al. (2024c), and TOVA Oren et al. (2024) simplify attention patterns by discarding KV caches mid-way through the context. However, these approaches struggle to retain the original models’ long-context capabilities due to limited global context modeling. Recent works like DuoAttention Xiao et al. (2024), RetrievalAttention Liu et al. (2024a), and SeerAttention Gao et al. (2024) address this issue by introducing retrieval heads Wu et al. (2024) or combining full attention with local attention heads. Quest Tang et al. (2024) introduces dynamic, query-aware sparsity for accelerated decoding, while MInference Jiang et al. (2024b) extends similar ideas to the prefilling stage. FastGen Ge et al. (2023) optimizes decoding by profiling attention heads to discard tokens. PQCache Zhang et al. (2024a) and ShadowKV Sun et al. (2024) further advance the selective attention methods with product quantization and low-rank decomposition. Additionally, LongLoRA Chen et al. (2023) finetunes short-context LLMs to long-context ones after converting global attention to shifted sparse attention.
@@ 7 Conclusion
We introduce LServe, an efficient serving system for long-sequence LLMs that leverages hybrid sparse attention. By incorporating unified block sparse attention, we achieve significant acceleration of the attention mechanism for both prefilling and decoding stages in long-sequence models. We further show that head-level static sparsity and query-aware dynamic sparsity are orthogonal and can be effectively combined with minimal impact on accuracy. LServe surpasses state-of-the-art systems, delivering an average of 1.3$\times$-2.1$\times$ speedup in the decoding stage and up to 2.9$\times$ speedup in the prefilling stage, preserving the models’ long-context capabilities.
@@ Acknowledgements
We thank MIT-IBM Watson AI Lab, MIT AI Hardware Program, MIT Amazon Science Hub, National Science Foundation, and Hyundai for supporting this research. We also thank June Yang, Bo Li, and Kaiyu Xie for their helpful discussions.

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable potential in processing long sequences and complex reasoning tasks, yet efficiently serving these models remains challenging due to the quadratic computational complexity of attention in the prefilling stage and the large memory footprint of the KV cache in the decoding stage. To address these issues, we introduce LServe, an efficient system that accelerates long-sequence LLM serving via hybrid sparse attention. This method unifies different hardware-friendly, structured sparsity patterns for both prefilling and decoding attention into a single framework, where computations on less important tokens are skipped block-wise. LServe demonstrates the compatibility of static and dynamic sparsity in long-context LLM attention. This design enables multiplicative speedups by combining these optimizations. Specifically, we convert half of the attention heads to nearly free streaming heads in both the prefilling and decoding stages. Additionally, we find that only a constant number of KV pages is required to preserve long-context and reasoning capabilities, irrespective of context length. We then design a hierarchical KV page selection policy that dynamically prunes KV pages based on query-centric similarity. On average, LServe accelerates LLM prefilling by up to 2.9$\times$ and decoding by 1.3-2.1$\times$ over vLLM, maintaining long-context accuracy. Code is released at https://github.com/mit-han-lab/omniserve.
footnotetext: *indicates equal contribution. Part of the work was done when Shang Yang and Haotian Tang were interning at NVIDIA. 1MIT 2SJTU 3NVIDIA. Correspondence to: Song Han <songhan@mit.edu>.
Proceedings of the $\mathit{8}^{th}$ MLSys Conference, Santa Clara, CA, USA, 2025. Copyright 2025 by the author(s).
@@ 1 Introduction
Large Language Models (LLMs) have dramatically transformed the field of artificial intelligence. With expanding context window lengths and test-time computing OpenAI (2024); Snell et al. (2024), LLMs now demonstrate remarkable performance across diverse long-sequence applications Gemini Team, Google (2024), including multi-turn conversations, long document analysis Zhang et al. (2024b); Goyal & Durrett (2020); Huang et al. (2021), multi-modal understanding Xue et al. (2024a); Liu et al. (2024b); Lin et al. (2024a), code completion Li et al. (2023); Lozhkov et al. (2024), and complex reasoning tasks  DeepSeek-AI et al. (2025). Many of these applications require processing hundreds of thousands of context tokens in real-world settings, presenting unique challenges. In particular, the demand for fast prefilling, or minimizing the time to the first token, and the burden on the decoding phase due to the large KV (key-value) caches necessary for such contexts, represent significant hurdles.
@@
We benchmark LServe across three long-sequence LLMs—Llama-3-8B, Minitron-4B, and Llama-2-7B—at context lengths up to 512k tokens. Compared to state-of-the-art frameworks like vLLM Kwon et al. (2023b), QServe Lin et al. (2024b), MInference Jiang et al. (2024b), and DuoAttention Xiao et al. (2024), LServe accelerates prefilling stage by up to 2.9$\times$ and achieves an average of 1.3$\times$-2.1$\times$ speedup in the decoding stage, without sacrificing the long-context capabilities of the original dense, floating-point models. Furthermore, LServe also demonstrates compatibility with the latest reasoning-centric models. Table 4 indicates that LServe achieves comparable accuracy to dense baselines on DeepSeek-R1-Distill-Llama-8B DeepSeek-AI et al. (2025).
@@ 2.2 Motivation
Figure 2: Latency breakdown of LLM inference during prefilling and decoding stages. Attention dominates both stages as sequence length increases, due to its quadratic complexity in prefilling and linear complexity in decoding. GEMM exhibits linear complexity in prefilling and constant complexity in decoding. Measurements obtained with Llama-3-8B on NVIDIA A100 GPU.
@@ 4.1 Evaluation Setup
Table 4: Accuracy evaluation on AIME MAA (2024) and MATH500 Hendrycks et al. (2021) DS-R1-Llama-8B refers to DeepSeek-R1-Distill-Llama-8B. DeepSeek-AI et al. (2025)
Model DS-R1-Llama-8B
Benchmark Dense LServe
AIME@2024 43.3 43.3
MATH500 84.2 85.4
Average 63.8 64.4
@@ 4.2 End-to-end Accuracy
Additionally, we also evaluate LServe on complex reasoning tasks such as AIME2024 MAA (2024) and MATH500 Hendrycks et al. (2021). As shown in Table 4, LServe maintains accuracy comparable to its dense counterparts on these complex reasoning tasks.
@@ 6 Related Work
Sparse Attention. BigBird Zaheer et al. (2020) reduces attention complexity by blending local, global, and random attention masks. Subsequent methods like StreamingLLM Xiao et al. (2023), H2O Zhang et al. (2024c), and TOVA Oren et al. (2024) simplify attention patterns by discarding KV caches mid-way through the context. However, these approaches struggle to retain the original models’ long-context capabilities due to limited global context modeling. Recent works like DuoAttention Xiao et al. (2024) and RetrievalAttention Liu et al. (2024a) address this issue by introducing retrieval heads Wu et al. (2024) or combining full attention with local attention heads. SeerAttention Gao et al. (2024) introduces a learnable gate to identify block-level attention sparsity. Quest Tang et al. (2024) applies dynamic, query-aware sparsity for accelerated decoding, while MInference Jiang et al. (2024b) extends similar ideas to the prefilling stage. FastGen Ge et al. (2023) optimizes decoding by profiling attention heads to discard tokens. PQCache Zhang et al. (2024a) and ShadowKV Sun et al. (2024) further advance the selective attention methods with product quantization and low-rank decomposition. Additionally, LongLoRA Chen et al. (2023) finetunes short-context LLMs to long-context ones after converting global attention to shifted sparse attention.
@@ 7 Conclusion
We introduce LServe, an efficient serving system for long-sequence LLMs that leverages hybrid sparse attention. By incorporating unified block sparse attention, we achieve significant acceleration of the attention mechanism for both prefilling and decoding stages in long-sequence models. We further show that head-level static sparsity and query-aware dynamic sparsity are orthogonal and can be effectively combined with minimal impact on accuracy. LServe surpasses state-of-the-art systems, delivering an average of 1.3$\times$-2.1$\times$ speedup in the decoding stage and up to 2.9$\times$ speedup in the prefilling stage, preserving the models’ capabilities to effectively handle long-context documents and perform complex reasoning tasks.
@@ Acknowledgements
We thank MIT-IBM Watson AI Lab, MIT AI Hardware Program, MIT Amazon Science Hub, and National Science Foundation for supporting this research. We also thank June Yang, Bo Li, and Kaiyu Xie for their helpful discussions.
Appendix A Artifact Appendix
A.1 Abstract
This artifact contains necessary scripts and dependencies to faithfully reproduce the crucial experiments presented in the paper. To successfully run the experiments, a host system with x86_64 CPUs is required, along with at least one A100 or L40S NVIDIA GPU. We also provide a pre-built docker image to simplify the environment setup process.
A.2 Artifact check-list (meta-information)
  • Program: Accuracy evaluation and efficiency benchmarking code for LServe; efficiency benchmarking code for baseline systems such as vLLM.
  • Compilation: Completed in the docker.
  • Transformations: N/A.
  • Binary: N/A.
  • Model: We provide a quantized version of Llama-3-8B to simplify the evaluation process.
  • Data set: Included in the docker image.
  • Run-time environment: NVIDIA Container Toolkit (nvidia-docker).
  • Hardware: A host with x86_64 CPUs and at least one NVIDIA A100 GPU (recommended) or L40S GPU.
  • Run-time state: N/A.
  • Execution: All benchmarks are executed on NVIDIA GPUs, while some data pre-processing code is executed on the host CPU;
  • Metrics: Long-context benchmarks accuracy; LLM generation throughput.
  • Output: Accuracy numbers and generation throughput (tokens/second).
  • Experiments: Inference speed measurement for LServe and baseline systems such as vLLM; accuracy evaluation of LServe.
  • How much disk space required (approximately)?: 128G.
  • How much time is needed to prepare workflow (approximately)?: Around 1 hour to pull docker images depending on the Internet connection and CPU performance.
  • How much time is needed to complete experiments (approximately)?: Around 2 hours to finish the efficiency benchmarks; and 1 hour to finish the accuracy benchmarks depending on the GPU performance and number of benchmark subsets to evaluate.
  • Publicly available?: Yes.
  • Code licenses (if publicly available)?: Apache License 2.0.
  • Data licenses (if publicly available)?: MIT.
  • Workflow framework used?: Docker.
  • Archived (provide DOI)?: https://doi.org/10.5281/zenodo.14989916
A.3 Description
A.3.1 How delivered
We will provide AE reviewers with a pre-built docker image containing LServe, vLLM, and all necessary dependencies.
A.3.2 Hardware dependencies
A host machine with x86_64 CPUs and at least one NVIDIA A100 GPU.
A.3.3 Software dependencies
A GPU-compatible Docker runtime environment is required.
A.3.4 Data sets
We require datasets such as 2wikimqa, dureader, and hotpotqa for LongBench evaluation. All datasets are publicly available and included in the docker image.
A.4 Installation
We recommend that users utilize our pre-built Docker images to set up the environment and run all experiments within the GPU-supported Docker container.
1docker run --gpus all -it --workdir /workspace/projects shang12138/lserve-mlsys25-ae
A.5 Experiment workflow
A.5.1 LServe Accuracy Evaluation
We provide push-button solution for evaluating LServe on LongBench.
1cd /workspace/projects/omniserve
2bash eval/scripts/LongBench/submit_longbench_dense.sh
3 # Evaluate the baseline accuracy
4bash eval/scripts/LongBench/submit_longbench_sparse.sh
5 # Evaluate LServe accuracy
6
7# The evaluation results can be found at ./eval/LongBench/pred/<model-name>/result.json
A.5.2 Throughput Benchmark
The generation throughputs of LServe and baseline system (i.e., vLLM) can be measured with the following commands.
1# LServe benchmark
2cd /workspace/projects/omniserve
3bash scripts/lserve_benchmark/launch.sh
4# Results in ./results.csv
5
6# vLLM benchmark
7cd /workspace/projects/vllm
8bash launch_server.sh # Start vLLM server
9# When the vLLM server has been launched,
10# start a new terminal in the same docker
11cd /workspace/projects/vllm
12bash run_vllm.sh # Launch evaluation
13# Results in ./results.csv
A.6 Evaluation and expected result
We provide reference numbers for evaluation results in this section. Please note that absolute speed measurements may vary slightly, even on identical GPU platforms, due to differences in machine conditions. However, the relative acceleration ratios should remain consistent. Additionally, due to randomness inherent in LLM generation, accuracy results might show minor deviations from the reported reference numbers.
Sequence Length vLLM (ms) LServe (ms) Speedup
64k 12.51 11.49 1.09x
96k 14.49 12.05 1.20x
128k 16.34 12.74 1.28x
160k 18.20 12.88 1.41x
192k 21.73 13.30 1.63x
224k 21.96 13.73 1.60x
256k 23.72 14.20 1.67x
320k 27.45 15.10 1.82x
Table 7: Generation latency of LServe and baseline (vLLM).
Model Llama-3-8B
Benchmark Dense LServe
2WikiMQA 26.2 27.0
DuReader 22.3 25.6
HotpotQA 41.1 40.8
MultiNews 27.6 27.1
Qasper 29.1 28.5
Table 8: LongBench evaluation results.
A.7 Methodology
Submission, reviewing and badging methodology: